# Asking Fast Radio Bursts (FRBs) for More than Reionization History

### v1

@@ 4 EoR Simulation Using grizzly
We now use the simulated EoR ionization field LCs to estimate the ${\rm DM}$ and other related quantities to examine how the different reionization models affect them. We simulated three EoR $x_{\rm{H}\,\textsc{ii}}$ LCs corresponding to the three different reionization histories $\bar{x}_{\rm{H}\,\textsc{ii}}$ as shown in Figure 9. In the following sections, we briefly describe the method used to simulate the EoR scenarios followed by all the corresponding estimates.
Figure 10: Light-cone slices of the ionization field. The three different panels correspond to the LC slices obtained from the realistic grizzly simulations for Slower, Fiducial and Faster reionization histories as labelled.
@@ 4.1 EoR Simulations
The EoR $x_{\rm{H}\,\textsc{ii}}$ LCs used here are constructed by stitching thin slices from the coeval $x_{\rm{H}\,\textsc{ii}}$ cubes simulated at several different redshifts in chronological order. We simulate $x_{\rm{H}\,\textsc{ii}}$ coeval boxes using a CD-EoR code, grizzly (Ghara et al., 2015a, 2018), which is based on a 1D radiative transfer technique. This algorithm takes the dark-matter density field and the corresponding halo catalogue from a $N$-body simulation to produce a $x_{\rm{H}\,\textsc{ii}}$ map at a redshift for a particular astrophysical source model.
@@
We use the dark-matter density fields and the associated halo catalogues obtained from the PRACE [^2] Partnership for Advanced Computing in Europe: http://www.prace-ri.eu/ project, PRACE4LOFAR. The input dark-matter distributions are generated using dark-matter only N-body simulation code cubep3m  [^3] http://wiki.cita.utoronto.ca/mediawiki/index.php/CubePM (Harnois-Déraps et al., 2013). The 3D density cubes have comoving volume $[500~{}h^{-1}~{}{\rm Mpc}]^{3}$ (see e.g, Dixon et al., 2015; Giri et al., 2019) which are gridded into $[600]^{3}$ voxels. The dark-matter particle distributions have been used to find the collapsed halos using spherical overdensity halo finder (Watson et al., 2013). The minimum halo mass in the PRACE4LOFAR simulation is $\sim 10^{9}\,{\rm M}_{\odot}$, which corresponds to $\approx 25$ dark-matter particles. We simulated $63$ coeval dark-matter cubes between a redshift range $6.1\lesssim z\lesssim 15.6$ with an equal time gap of $11.4~{}{\rm Myr}$.
@@
We consider an EoR source model where the dark-matter halos with masses larger than $10^{9}~{}{\rm M}_{\odot}$ host UV photon-emitting galaxies. We assume that the stellar mass of a galaxy ($M_{\star}$) is related to the host dark matter halo mass $M_{\rm halo}$ as $M_{\star}\propto M^{\alpha_{s}}_{\rm halo}$. We tune the ionization efficiency ($\zeta$) so that the reionization ends at $z\sim 6$. Note that all reionization models considered in this study are inside-out in nature. Our fiducial grizzly model corresponds to a choice of $\alpha_{s}=1$ and spans from redshift $6$ to $15.6$. We consider a rapidly evolving Faster and a slowly evolving Slower reionization scenario which correspond to $\alpha_{s}=2$ and $\alpha_{s}=0.1$ respectively.
@@
Next, we produce coeval cubes of $x_{\rm{H}\,\textsc{ii}}$ at $63$ redshifts between $6.1$ and $15.6$ for the aforementioned source model. We refer the readers to Ghara et al. (2015a) and Islam et al. (2019) for more details about these calculations. Finally, we used these coeval cubes of $x_{\rm{H}\,\textsc{ii}}$ to create the LC which accounts for the evolution of $x_{\rm{H}\,\textsc{ii}}$ with redshift. The detailed method to implement the LC effect can be found in Ghara et al. (2015b). The reionization histories for the three different EoR scenarios are shown in Figure 9 while we present the corresponding LCs in Figure 10. The LCs clearly show the difference in the patchiness of the reionization scenarios.
@@
In Figure 11, we show the ${\rm\overline{{DM}}_{EoR}}$ estimated from the LC simulations for ‘Slower’, ‘Fiducial’ and ‘Faster’ reionization histories. The solid lines show the mean estimated using all the LoS (here $360~{}000$ grid points) in the simulations and the shaded regions around them are the respective $1\sigma$ deviations due to cosmic variance. We consider small $z$-bins ($\Delta z=0.5$) while computing the mean and the sample variance. We overplot the ${\rm\overline{{DM}}_{EoR}}$ estimates calculated using the average ionization fraction (Figure 9) for the three reionization scenarios.
@@
${\rm\overline{{DM}}_{EoR}}(z)$ begins with a near zero value (with low-$z$ IGM contributions subtracted) around the end of reionization ($z\approx 6$) and increases almost linearly with $z$ until when ionization is sufficiently small ($\bar{x}_{\rm{H}\,\textsc{ii}}\approx 0.1$), where it starts to plateau. This is qualitatively consistent across both the realistic and toy model reionization histories (see Figure 2). The rate of linear rise at lower $z$ and the saturation value is highest for the slower reionization history, and these decrease monotonically for faster histories. The ‘knee’ in the ${\rm\overline{{DM}}_{EoR}}(z)$ curves occurs at a larger redshift for slower reionization histories. The saturation values of ${\rm DM}$ differ by more than $1\sigma$ across all three histories.
@@
Figure 11: Sky-averaged DM and the corresponding cosmic variance during EoR. The solid lines represent the ${\rm\overline{{DM}}_{EoR}}$ estimates computed from realistic grizzly LC simulations. Shaded regions around them are the corresponding $1\sigma$ cosmic variance. The cyan, green and orange colors respectively represent Slower, Fiducial and Faster reionization histories. The three dashed lines represent the ${\rm\overline{{DM}}_{EoR}}$ computed using the globally-averaged ionization fractions $\bar{x}_{\rm{H}\,\textsc{ii}}$ shown in Figure 9.
@@
The ${\rm\overline{{DM}}_{EoR}}(z)$ and ${\rm DM}(\bar{x}_{\rm{H}\,\textsc{ii}}(z))$ are significantly distinct for all three reionization histories due to IGM electron density fluctuations driven by the formation and growth of ionized bubbles, as well as the underlying density perturbations. The bias due to the fluctuations are highest at larger $z$ ($z\gtrsim 12$) where the ${\rm DM}$ saturates, and the differences between the two estimates decrease rapidly towards the end of reionization. This is because the merging and percolation of the ionized bubbles typically wash out the fluctuations towards the mid and advanced stages of reionization. The difference is smallest ($\lesssim 1\sigma$) for the Faster reionization scenario, where large ionization bubbles appear suddenly at lower redshifts (see bottom panel of Figure 10) and quickly percolate and fill up the entire IGM. In this case, the IGM is more patchy and hence small-scale fluctuations do not prevail. Relatively small ionized bubbles form for slower histories and takes a longer time to fill the entire IGM. This leads to relatively more small-scale fluctuations (see Figure 10), thereby resulting in a larger difference ($\gtrsim 2\sigma$) between the ${\rm\overline{{DM}}_{EoR}}(z)$ and ${\rm DM}(\bar{x}_{\rm{H}\,\textsc{ii}}(z))$ (solid and dashed lines) for the slower reionization history.
@@
Figure 12: Derivative of ${\rm\overline{{DM}}_{EoR}}(z)$ w.r.t. redshift $z$. The three lines show the derivatives obtained from realistic simulations shown in Figure 10.
As the average ${\rm\overline{{DM}}_{EoR}}$ is an integrated quantity, we compute its derivative $d{{\rm\overline{{DM}}_{EoR}}}/dz$ to extract instantaneous information at a given stage of reionization. We show the derivatives in Figure 12 for the three histories. $d{\rm\overline{{DM}}_{EoR}}/dz$ starts with a high value which is similar for the three histories. This simply indicates that the electron distribution in the IGM is roughly the same for all three scenarios nearing the end of reionization. The large values of the derivative at $z\lesssim 8$ are an implication of the fact that $\bar{x}_{\rm{H}\,\textsc{ii}}$ changes rapidly during the mid and end stages of the reionization. However, the difference in how fast $\bar{x}_{\rm{H}\,\textsc{ii}}(z)$ evolves causes the derivatives to distinctly vary for the three histories considered here. The derivative for the faster scenario has small values but is the steepest among the three. The value of the derivatives increases for the fiducial and further to slower reionization histories, while the slopes of the curves decreases.
Figure 13 depicts $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ as a function of $z$. This shows how the sample variance error in the measurement is related with the mean DM. The qualitative behavior of the ratio is similar for the three reionization histories. The ‘knee’ of saturation in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ depends on the $z_{\rm mid}$ as seen in toy models (top panel of Figure 3). However, contrary to the toy models (Figures 3 and 6), the ratio $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ decreases with $z$. The $1\sigma$ cosmic variance for all the three histories are roughly similar, hence the ratio scales inversely with ${\rm\overline{{DM}}_{EoR}}$. This could be because of the fact that in the real simulations the electron distribution in the IGM has an additional contribution coming from the underlying matter density field which is absent in our toy models. This needs further detailed investigation with a more sophisticated toy model and we defer it to a future work.
@@
Figure 14 shows contours of the derivatives of the structure function $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ on the $\delta\theta-z$ plane, similar to Figures 4 and 7. We estimate $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$ from $400$ LoS per comoving slice and bin them within bins of $\Delta z=0.5$. The behaviour is qualitatively similar to the toy models. The contours are tightly squeezed towards lower $z$ ($10\%$ contour at $z\lesssim 7.2$) for the faster reionization history and they gradually expand for the fiducial and slower models ($10\%$ contour at $z\lesssim 8.1$). The effects of the different IGM topologies are also clearly evident from the $\delta\theta$ extent of the contours. The faster history has larger bubbles (hence patchy IGM) resulting in extended angular correlations. In contrast, the angular correlations are smaller (depicted by the squeezed contours along $\delta\theta$) for slower histories where the bubbles are smaller in sizes. This trend closely resembles what is observed in Figure 7.
@@
Figure 14: Density distribution of $\partial^{2}\Xi/\partial z~{}\partial\delta\theta$. This plot is similar to Figure 4, but for the realistic EoR simulations. The three different panels correspond to the three different reionization histories as labelled. However, this is obtained from a realistic simulation of EoR LCs which does not have a fixed bubble size.
4.3 Impact of Non-uniform Distribution of FRBs and Post-EoR
@@
All the analyses presented above are independent of the distribution of FRBs in a LC box. There are various factors that make the distribution of FRBs non-uniform in the sky-plane and across redshift. We repeat the analysis for the grizzly simulated LCs considering a realistic redshift distribution of the sky-averaged number of FRBs. Following the estimates in Beniamini et al. (2021), we used the cumulative observable redshift distribution as shown in Figure 15. To make our estimates more realistic, we also consider the FRBs to be correlated with the overdensities in our matter density LC. This is well motivated because the overdense regions are the places which can host radiation sources and therefore also FRBs. Ideally, one should correlate the FRBs with the halo field which is precisely the location of sources. However, here we lack the exact halo locations and instead use the matter overdensity peaks as a proxy. We denote by $N_{\rm FRB}^{\rm tot}$ the total number of FRBs observed in the range $6\leq z\leq 15$. We divide the whole EoR redshift range ($z=6$ to $15$) into bins of $\Delta z=0.5$ and estimate the number of FRBs per bin according to the distribution (Figure 15) for a given $N_{\rm FRB}^{\rm tot}$. For each bin, we randomly pick up the grid points which are biased by the high-dense regions. We finally use those grid points to estimate the average ${\rm DM}$ and corresponding dispersion for all the redshift bins.
Figure 16: Variation of $\langle\overline{{\rm DM}}(z_{\rm mid})\rangle$ and the marginalized $\langle\overline{\rm DM}_{\rm margin}\rangle$ as a function of total number of FRBs $N_{\rm FRB}^{\rm tot}$ observed during the EoR. Top row: This shows the ensemble averaged $\overline{{\rm DM}}(z_{\rm mid})$ and its standard deviation as a function of $N^{\rm tot}_{\rm FRB}$ for the three reionization histories. Bottom row: This shows ensemble mean of $\overline{\rm DM}_{\rm margin}$, marginalized $\overline{\rm DM}$ over the EoR redshift range, as a function of $N_{\rm FRB}^{\rm tot}$ for the three different reionization histories. Solid lines here are the ensemble mean and shaded regions correspond to $3\sigma$ cosmic variances which are estimated, for every choice of $N_{\rm FRB}^{\rm tot}$, using $1000$ independent ensembles of FRB distributions in the IGM during EoR. The left, middle and right columns correspond to the optimistic, moderate and pessimistic scenarios, respectively, as described in the text.

### v2

@@ 4 EoR Simulation Using grizzly
We now use the simulated EoR ionization field LCs to estimate the $\rm DM$} and other related quantities to examine how the different reionization models affect them. We simulated three EoR $x_{\rm{H}\,\textsc{ii}}$ LCs corresponding to the three different reionization histories $\bar{x}_{\rm{H}\,\textsc{ii}}$ as shown in Figure 9. In the following sections (§4.1-§4.3), we briefly describe the method used to simulate the EoR scenarios, followed by all the corresponding estimates.
Figure 10: Light-cone slices of the ionization field. The three different panels correspond to the LC slices obtained from the realistic grizzly simulations for Slower, Fiducial, and Faster reionization histories, as labeled.
@@ 4.1 EoR Simulations
The EoR $x_{\rm{H}\,\textsc{ii}}$ LCs used here are constructed by stitching thin slices from the coeval $x_{\rm{H}\,\textsc{ii}}$ cubes simulated at several different redshifts in chronological order. We simulate $x_{\rm{H}\,\textsc{ii}}$ coeval boxes using an EoR code, grizzly (Ghara et al., 2015a, 2018), which is based on a 1D radiative transfer technique. This algorithm takes the dark-matter density field and the corresponding halo catalog from an $N$-body simulation to produce a $x_{\rm{H}\,\textsc{ii}}$ map at a redshift for a particular astrophysical source model.
@@
We use the dark-matter density fields and the associated halo catalogs obtained from the PRACE [^3] Partnership for Advanced Computing in Europe: http://www.prace-ri.eu/ project, PRACE4LOFAR. The input dark-matter distributions are generated using dark matter only N-body simulation code cubep3m  [^4] https://wiki.cita.utoronto.ca/index.php/CubePM (Harnois-Déraps et al., 2013). The 3D density cubes have comoving volume $500~h^{-1}~{\rm Mpc}]^{3}$ (see, e.g, Dixon et al., 2015; Giri et al., 2019) which are gridded into $600]^{3}$ voxels. The dark-matter particle distributions have been used to find the collapsed halos using spherical overdensity halo finder (Watson et al., 2013). The minimum halo mass in the PRACE4LOFAR simulation is $\sim 10^{9}\,{M}_{\odot}$, which corresponds to $\approx 25$ dark-matter particles. We simulated $63$ coeval dark-matter cubes between a redshift range $6.1\lesssim z\lesssim 15.6$ with an equal time gap of $11.4~{\rm Myr}$.
@@
We consider an EoR source model where the dark-matter halos with masses larger than $10^{9}~{M}_{\odot}$ host UV photon-emitting galaxies. We assume that the stellar mass of a galaxy ($M_{\star}$) is related to the host dark-matter halo mass $M_{\rm halo}$ as $M_{\star}\propto M^{\alpha_{s}}_{\rm halo}$. We tune the ionization efficiency [^5] The ionization efficiency is kept constant for a given simulation and assumed to be independent of $z$. ($\zeta$) so that the reionization ends at $z\sim 6$. Note that all reionization models considered in this study are inside-out in nature. Our fiducial grizzly model corresponds to a choice of $\alpha_{s}=1$ and spans from redshift $6$ to $15.6$. We consider a rapidly evolving Faster and a slowly evolving Slower reionization scenario, which correspond to $\alpha_{s}=2$ and $\alpha_{s}=0.1$, respectively. Here, $\alpha_{s}=1+\alpha_{\star}+\alpha_{\rm esc}$, with $\alpha_{\star}$ and $\alpha_{\rm esc}$ representing the power-law indices of the halo mass dependence for the star formation rate and the UV photon escape fraction, respectively (e.g., Park et al., 2019). A typical range of $\alpha_{\star}\in[-0.5,1]$ and $\alpha_{\rm esc}\in[-1,0.5]$ leads to $\alpha_{s}\in[-0.5,2.5]$.
@@
Next, we produce coeval cubes of $x_{\rm{H}\,\textsc{ii}}$ at $63$ redshifts between $6.1$ and $15.6$ for the aforementioned source model. We refer the readers to Ghara et al. (2015a) and Islam et al. (2019) for more details about these calculations. Finally, we used these coeval cubes of $x_{\rm{H}\,\textsc{ii}}$ to create the LC, which accounts for the evolution of $x_{\rm{H}\,\textsc{ii}}$ with redshift. The detailed method to implement the LC effect can be found in Ghara et al. (2015b). The reionization histories for the three different EoR scenarios are shown in Figure 9 while we present the corresponding LCs in Figure 10. The LCs clearly show the difference in the patchiness of the reionization scenarios.
@@
In Figure 11, we show the $\rm\overline{{DM}}_{EoR}$} estimated from the LC simulations for ‘Slower,’ ‘Fiducial,’ and ‘Faster’ reionization histories. The solid lines show the mean estimated using all the LoS (here $360,000$ grid points) in the simulations, and the shaded regions around them are the respective $1\sigma$ deviations due to cosmic variance. We consider small $z$-bins ($\Delta z=0.5$) while computing the mean and the sample variance. We overplot the $\rm\overline{{DM}}_{EoR}$} estimates calculated using the average ionization fraction (Figure 9) for the three reionization scenarios.
@@
${\rm\overline{{DM}}_{EoR}}(z)$ begins with a near-zero value (with low-$z$ IGM contributions subtracted) around the end of reionization ($z\approx 6$) and increases almost linearly with $z$ until when ionization is sufficiently small ($\bar{x}_{\rm{H}\,\textsc{ii}}\approx 0.1$), where it starts to plateau. This is qualitatively consistent across both the realistic and toy model reionization histories (see Figure 2). The rate of linear rise at lower $z$ and the saturation value are highest for the slower reionization history, and these decrease monotonically for faster histories. The ‘knee’ in the ${\rm\overline{{DM}}_{EoR}}(z)$ curves occurs at a larger redshift for slower reionization histories. The saturation values of $\rm DM$} differ by more than $1\sigma$ across all three histories.
@@
Figure 11: Sky-averaged DM and the corresponding cosmic variance during EoR. The solid lines represent the $\rm\overline{{DM}}_{EoR}$} estimates computed from realistic grizzly LC simulations. Shaded regions around them are the corresponding $1\sigma$ cosmic variance. The cyan, green, and orange colors, respectively represent Slower, Fiducial, and Faster reionization histories. The three dashed lines represent the $\rm\overline{{DM}}_{EoR}$} computed using the globally averaged ionization fractions $\bar{x}_{\rm{H}\,\textsc{ii}}$ shown in Figure 9.
@@
The ${\rm\overline{{DM}}_{EoR}}(z)$ and ${\rm DM}(\bar{x}_{\rm{H}\,\textsc{ii}}(z))$ are significantly distinct for all three reionization histories due to IGM electron density fluctuations driven by the formation and growth of ionized bubbles, as well as the underlying density perturbations. The difference between the ${\rm\overline{{DM}}_{EoR}}(z)$ and ${\rm DM}(\bar{x}_{\rm{H}\,\textsc{ii}}(z))$ was not significant for toy models in Fig. 2 since the contribution to fluctuations due to underlying density perturbations $\Delta(\bm{\theta},z)$ is absent in our toy model. The bias due to the fluctuations is highest at larger $z$ ($z\gtrsim 12$) where the $\rm DM$} saturates, and the differences between the two estimates decrease rapidly toward the end of reionization. This is because the merging and percolation of the ionized bubbles typically wash out the fluctuations toward the mid and advanced stages of reionization. The difference is smallest ($\lesssim 1\sigma$) for the Faster reionization scenario, where large ionization bubbles appear suddenly at lower redshifts (see bottom panel of Figure 10) and quickly percolate and fill up the entire IGM. In this case, the IGM is more patchy and hence small-scale fluctuations do not prevail. Relatively small ionized bubbles form for slower histories and take a longer time to fill the entire IGM. This leads to relatively more small-scale fluctuations (see Figure 10), thereby resulting in a larger difference ($\gtrsim 2\sigma$) between the ${\rm\overline{{DM}}_{EoR}}(z)$ and ${\rm DM}(\bar{x}_{\rm{H}\,\textsc{ii}}(z))$ (solid and dashed lines) for the slower reionization history.
@@
Figure 12: Derivative of ${\rm\overline{{DM}}_{EoR}}(z)$ with respect to redshift $z$. The three lines show the derivatives obtained from realistic simulations shown in Figure 10.
@@
As the average $\rm\overline{{DM}}_{EoR}$} is an integrated quantity, we compute its derivative $d{{\rm\overline{{DM}}_{EoR}}}/dz$ to extract instantaneous information at a given stage of reionization. We show the derivatives in Figure 12 for the three histories. $d{\rm\overline{{DM}}_{EoR}}/dz$ starts with a high value, which is similar for the three histories. This simply indicates that the electron distribution in the IGM is roughly the same for all three scenarios nearing the end of reionization. The large values of the derivative at $z\lesssim 8$ are an implication of the fact that $\bar{x}_{\rm{H}\,\textsc{ii}}$ changes rapidly during the mid and end stages of the reionization. However, the difference in how fast $\bar{x}_{\rm{H}\,\textsc{ii}}(z)$ evolves causes the derivatives to distinctly vary for the three histories considered here. The derivative for the faster scenario has small values but is the steepest among the three. The value of the derivatives increases for the fiducial and further to slower reionization histories, while the slopes of the curves decrease. Comparing with the toy models (Fig. 3), we find the derivatives to have similar values, roughly for $z<10$. However, $d{{\rm\overline{{DM}}_{EoR}}}/dz$ is an order of magnitude smaller for simulation than in the toy model for $z>10$. The derivative is strongly dependent on the progress of the reionization history. The reason is mainly due to the difference in how fast reionization progresses for the toy models and simulations at those redshifts. The $\bar{x}_{\rm{H}\,\textsc{ii}}$ almost saturates in simulations for $z>10$, causing a reduced value of the derivatives.
@@
Figure 14: Density distribution of $\partial^{2}\Xi/\partial z~\partial\delta\theta$. This plot is similar to Figure 4, but for the realistic EoR simulations. The three different panels correspond to the three different reionization histories as labeled. However, this is obtained from a realistic simulation of EoR LCs, which does not have a fixed bubble size.
Figure 13 depicts $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ as a function of $z$. This shows how the sample variance error in the measurement is related to the mean DM. The qualitative behavior of the ratio is similar for the three reionization histories. The ‘knee’ of saturation in $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ depends on the $z_{\rm mid}$ as seen in toy models (top panel of Figure 3). However, contrary to the toy models (Figures 3 and 6), the ratio $\sigma_{\rm DM}/{\rm\overline{{DM}}_{EoR}}$ decreases with $z$. The $1\sigma$ cosmic variances for all three histories are roughly similar, hence the ratio scales inversely with $\rm\overline{{DM}}_{EoR}$}. This could be because of the fact that in the real simulations, the electron distribution in the IGM has an additional contribution coming from the underlying matter density field, which is absent in our toy models. This needs further detailed investigation with a more sophisticated toy model, and we defer it to a future work.
Figure 14 shows contours of the derivatives of the structure function $\partial^{2}\Xi/\partial z~\partial\delta\theta$ on the $\delta\theta-z$ plane, similar to Figures 4 and 7. We estimate $\partial^{2}\Xi/\partial z~\partial\delta\theta$ from $400$ LoS per comoving slice and bin them within bins of $\Delta z=0.5$. Although the contours look different from those in Figs. 4 and 7, the behavior is qualitatively similar to the toy models. The contours are tightly squeezed toward lower $z$ ($10\%$ contour at $z\lesssim 7.2$) for the faster reionization history, and they gradually expand for the fiducial and slower models ($10\%$ contour at $z\lesssim 8.1$). The effects of the different IGM topologies are also clearly evident from the $\delta\theta$ extent of the contours. The faster history has larger bubbles (hence patchy IGM), resulting in extended angular correlations. In contrast, the angular correlations are smaller (depicted by the squeezed contours along $\delta\theta$) for slower history where the bubbles are smaller in size. This trend closely resembles what is observed in Figure 7. The contours here are open along $\delta\theta$ as compared to the toy models. The reason is the difference in the morphology of the ionized regions in the toy model and the realistic simulations. Being a binary field created using spherical bubbles, the ionized regions in toy models have very sharp boundaries showing an abrupt change in free electron distribution (a step function; zero within the ionized regions and one outside). However, in simulations, the IGM can be partially ionized, which is a result of the local inhomogeneity in the IGM due to perturbations in the density and photon fields. This gives rise to a more complex morphology and ill-defined geometric shapes of the ionized regions in realistic simulations, resulting in a mixture of bubble scales in the IGM. Therefore, in simulations, we can visualize it in a way where contours of several bubble sizes are superposing and giving rise to the open contours and insignificant squeezing along $z$-axis.
4.3 Impact of Nonuniform Distribution of FRBs and Post-EoR
All the analyses presented above are independent of the distribution of FRBs in an LC box. There are various factors that make the distribution of FRBs nonuniform in the sky plane and across redshift. We repeat the analysis for the grizzly simulated LCs, considering a realistic redshift distribution of the sky-averaged number of FRBs. Following the estimates in Beniamini et al. (2021), we used the cumulative observable redshift distribution as shown in Figure 15. To make our estimates more realistic, we also consider the FRBs to be correlated with the overdensities in our matter density LC. This is well motivated because the overdense regions are the places that can host radiation sources and therefore also FRBs. Ideally, one should correlate the FRBs with the halo field, which is precisely the location of sources. However, here we lack the exact halo locations and instead use the matter overdensity peaks as a proxy. We denote by $N_{\rm FRB}^{\rm tot}$ the total number of FRBs observed in the range $6\leq z\leq 15$. We divide the whole EoR redshift range (z=615z=6-15) into bins of $\Delta z=0.5$ and estimate the number of FRBs per bin according to the distribution (Figure 15) for a given $N_{\rm FRB}^{\rm tot}$. For each bin, we randomly pick the grid points that are biased by the high-density regions. We finally use those grid points to estimate the average $\rm DM$} and corresponding dispersion for all the redshift bins.
@@ 4.3 Impact of Nonuniform Distribution of FRBs and Post-EoR
Figure 16: Variation of $\langle\overline{{\rm DM}}(z_{\rm mid})\rangle$ and the marginalized $\langle\overline{\rm DM}_{\rm margin}\rangle$ as a function of total number of FRBs $N_{\rm FRB}^{\rm tot}$ observed during the EoR. Top row: this shows the ensemble-averaged $\overline{{\rm DM}}(z_{\rm mid})$ and its standard deviation as a function of $N^{\rm tot}_{\rm FRB}$ for the three reionization histories. Bottom row: this shows ensemble mean of $\overline{\rm DM}_{\rm margin}$, marginalized $\overline{\rm DM}$ over the EoR redshift range, as a function of $N_{\rm FRB}^{\rm tot}$ for the three different reionization histories. Solid lines here are the ensemble mean, and shaded regions correspond to $3\sigma$ cosmic variances, which are estimated, for every choice of $N_{\rm FRB}^{\rm tot}$, using $1000$ independent ensembles of FRB distributions in the IGM during EoR. The left, middle, and right columns correspond to the optimistic, moderate, and pessimistic scenarios, respectively, as described in the text.
@@
$\overline{\rm DM}_{\rm margin}=\frac{\int_{z_{\rm min}}^{z_{\rm max}}{\rm DM}_{\rm EoR}(z)~dP/dz~dz}{\int_{z_{\rm min}}^{z_{\rm max}}dP/dz~dz}~,$ (5)
We introduce the contribution from post-EoR ($0\leq z\leq 6$) and consider three different scenarios – (i) Optimistic, where the contribution from the post-EoR is perfectly modeled and completely removed from the data, (ii) Moderate, where the post-EoR contribution is imperfectly removed from the individual ${\rm DM}(z)$ such that the residual [^6] We computed the $\sigma_{\rm DM}$ for ${\rm DM}_{\rm IGM}(z=6)$ slice using different LoS from low-$z$ simulations of IGM. We generated a residual ${\rm DM}_{\rm IGM}$ field with mean zero (corresponding to the Optimistic scenario) and Gaussian scale of $0.5\sigma_{\rm DM}$, and added this to the ${\rm DM}_{\rm EoR}$. low-$z$ contribution remains within $0.5\sigma_{\rm DM}$ at $z=6$, and (iii) Pessimistic, where the post-EoR contribution is completely present for every FRB measured. Note that these three scenarios are differentiated based on how accurately we know the ${\rm DM}(\bm{\theta},z<6)$ for every observed LoS $\bm{\theta}$ above $z=6$. The assumption that we either know the redshifts of every FRB or at least can separate EoR FRBs from the post-EoR FRBs confidently is implicit within all three scenarios and is itself nontrivial. Apart from the fact that the boundary between EoR and post-EoR is still fuzzy, the erroneous/indeterminate FRB redshifts can cause a spill over the boundary. This may bias our estimators, and we plan to investigate this in greater detail in future studies.
We simulate a matter density LC within the redshift range $0\leq z\leq 6$ (comoving distance $8542~{\rm Mpc}$), and assume that the electrons linearly follow the underlying density contrast. We use the publicly available gadget4777https://wwwmpa.mpa-garching.mpg.de/gadget4/ (Springel et al., 2021) to simulate the dark-matter density fields within comoving cubes of volume $500~h^{-1}~{\rm Mpc}]^{3}$ at the lower redshifts. We choose the same cosmological parameters as mentioned at the end of §1. We also set the resolution (for Particle-Mesh) of the box to match those in the grizzly simulations. We simulated the particle distribution at $66$ redshifts within the range $0\leq z\leq 7$, roughly equidistant by $200~{\rm Myr}$. We generated the density coeval boxes on $600^{3}$ grids and then used them to make the density LC. We use a weighted linear interpolation scheme to interpolate the density fields at the desired redshift slices from the redshifts at which the coeval boxes are generated. We cannot simulate a large box spanning the whole range up to $z\leq 6$. Hence, we need to repeat the boxes along the LoS ($z$-axis) while creating the final LC.
The contribution from low redshifts increases both $\rm\overline{{DM}}_{EoR}$} and $\sigma_{\rm DM}(z)$, making the separation between the two different histories more challenging. In Figure 16, $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ for the three different histories start at different $N_{\rm FRB}^{\rm tot}$. This is because, for the slower history, $z_{\rm mid}$ is larger than that of the other histories, requiring a higher $N_{\rm FRB}^{\rm tot}$ to populate the corresponding $z$-bin with at least two FRBs. $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ converges very quickly for $N_{\rm FRB}^{\rm tot}\approx 10$; however, the corresponding dispersion is large, and it decreases with increasing $N^{\rm tot}_{\rm FRB}$, as expected. $N_{\rm FRB}^{\rm tot}\geq 20$ is sufficient to distinguish $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ at $3\sigma$ for the optimistic scenario. This lower bound remains the same to distinguish the slower reionization history from the faster one for the moderate scenario. However, it takes nearly $N_{\rm FRB}^{\rm tot}=150$ and $40$ to discern $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ of slower-fiducial and fiducial-faster pairs of histories, respectively. Despite having the same $\langle\overline{\rm DM}(z_{\rm mid})\rangle$, we need more FRBs due to additional fluctuations (with zero mean) coming from the lower $z$ in the moderate scenario. Considering the pessimistic case, it takes $N_{\rm FRB}^{\rm tot}=80,~150,~600$ to distinguish between faster-slower, slower-fiducial, and fiducial-faster pairs of histories, respectively. Since the low-$z$ contribution is being completely included in the pessimistic scenario, the values of $\langle\overline{\rm DM}(z_{\rm mid})\rangle$ increase in comparison with the other scenarios.
Figure 16 also shows the variation of the $\langle\overline{{\rm DM}}_{\rm margin}\rangle$ with $N_{\rm FRB}^{\rm tot}$. Considering the optimistic case, we need roughly $40$ FRBs identified during the EoR to discern slower and faster reionization histories at $3\sigma$ with $\langle\overline{{\rm DM}}_{\rm margin}\rangle$. Whereas the $N_{\rm FRB}^{\rm tot}$ slightly increases to $60$ and $90$ for faster-fiducial and fiducial-slower pairs of histories, respectively. These numbers respectively increase to $90$, $200$, and $300$ for the moderate case, as shown in the middle panel. Finally, for the pessimistic scenario, we need $N_{\rm FRB}^{\rm tot}\approx 220,~600,~1000$ to discern (at $3\sigma$) slower-faster, faster-fiducial, and fiducial-slower pairs of reionization histories, respectively. The numbers we found are realistic, and it would be possible to detect many FRBs during EoR using the upcoming SKA-Mid.
5 Discussion